OpenStudy (laacmichael):

If the equation y=1/6​​ (x+12) were graphed in the xy plane, which of the following statements would be true of the graphed line? 1 It would be perpendicular to the graph of y=1/6(x+12) 2 It would be parallel to the graph of 12y=2x+3 3 It would have the same slope as the graph of x+6y=18x 4 It would have the same yyy‑intercept as the graph of y=1/6x+12

OpenStudy (tkhunny):

Just put them all in the same form and compare. In this case, it seems to be asking about slopes and intercepts. Why not put them all in slope-intercept form? It's just a little algebra.

OpenStudy (laacmichael):

OK

OpenStudy (laacmichael):

@tkhunny 1 Y=1/6X+2 2 y=1/6x+1/4 3 y=17/6x 4 y=1/6x+12

OpenStudy (laacmichael):

So the awnser is 4 y=1/6x+12